# M4olGen: Multi-Agent, Multi-Stage Molecular Generation under Precise Multi-Property Constraints

### v1

@@ 2 Related Work
Deep generative models have been widely applied to molecular design, leveraging graph or sequence-based representations such as SMILES. Early works include VAEs (Gómez-Bombarelli et al., 2016) and GANs such as MolGAN (Cao and Kipf, 2018), followed by graph-based models like GCPN (You et al., 2018), GraphAF (Shi et al., 2020), and MoFlow (Zang and Wang, 2020). Spanning Tree Graph Generation (STGG) (Ahn et al., 2021; Jolicoeur-Martineau et al., 2025) shows promising performance in multi-objective optimization by combining sequential modeling with structured tree-based molecule representation. Reinforcement learning approaches (e.g., MolDQN (Zhou et al., 2018) and GFlowNets (Bengio et al., 2023; Jain et al., 2023)) enable property-driven optimization, often with multi-objective extensions for QED, LogP, and SA. However, these single-agent methods struggle to exactly satisfy multiple numeric constraints, reflecting exploration–exploitation trade-offs.
@@ 4.1 Experimental Setup
Table 1: Overall metrics across methods (lower is better). Best per column in bold; second best underlined.
@@
Table 2: Performance comparison across methods on HOMO/LUMO–constrained molecular generation. Lower is better for error metrics; higher is better for diversity and uniqueness.
@@ 4.2 Results and Analysis
Protocol. GRPO is ground-truth–free and reward-based, so performance is not tied to a particular training distribution. To test generalization, we uniformly sample 100 target tuples $(\mathrm{QED},\mathrm{LogP},\mathrm{MW})$ and $\mathrm{HOMO},\mathrm{LUMO}$ across admissible ranges. For each tuple and each baseline we run 10 independent trials under the same compute budget and report the best-of-10. Across settings, our normalized total error (NTE) decreases monotonically with hop count shown in both Table 1 and Table 2.
@@ 4.3 Ablation Study
Interpretation. We ablate three design choices on a held-out set: (i) Stage I without retrieval (baseline), (ii) Stage 1 with retrieval, and (iii) Stage I with retrieval followed by a fragment-level optimizer using 1/2/3 hops. We report per-property errors (QED, logP, MW) and the normalized total error ($e_{\mathrm{norm}}=\lvert\Delta\mathrm{QED}\rvert+\lvert\Delta\log P\rvert/10+\lvert\Delta\mathrm{MW}\rvert/700$) in Table 3. For visualization, we plot the drop percentage relative to the no-retrieval baseline,
@@
Effect of retrieval Adding retrieval already yields consistent gains: the normalized total error drops by 13.7% ($0.307\to 0.265$), driven primarily by improvements in logP ($\mathbf{20.7\%}$ drop) and MW ($7.8\%$ drop). Retrieval also gives the best stand-alone QED error among non-optimized variants ($\mathbf{0.098}$, $\mathbf{11.7\%}$ drop).
@@
Effect of the fragment-level optimizer Introducing the optimizer produces the largest improvements, especially on MW. Moving from retrieval-only to 1/2/3 hops reduces MW error from $\sim 63$ to $\sim 10$ ($\mathbf{84.9\%}$ drop vs. baseline), and steadily improves logP (drops of $56.4\%$, $65.1\%$, and $\mathbf{70.7\%}$). The overall normalized error decreases monotonically with more hops: $0.187$ (1-hop, $39.1\%$ drop), $0.160$ (2-hop, $47.9\%$ drop), and $\mathbf{0.146}$ (3-hop, $\mathbf{52.4\%}$ drop). QED exhibits a small regression at 1-hop (as expected when trading off multi-objective targets), but recovers by 3-hop to a $7.0\%$ drop versus baseline.
Takeaway Retrieval is a strong enabler, and the fragment-level optimizer is essential for precise multi-property alignment, culminating in the best overall performance with the 3-hop setting.

### v2

@@ 2 Related Work
Deep generative models have been widely applied to molecular design, leveraging graph or sequence-based representations such as SMILES. Early works include VAEs (Gómez-Bombarelli et al., 2016) and GANs such as MolGAN (Cao and Kipf, 2018), followed by graph-based models like GCPN (You et al., 2018), GraphAF (Shi et al., 2020), and MoFlow (Zang and Wang, 2020). Spanning Tree Graph Generation (STGG) (Ahn et al., 2021; Jolicoeur-Martineau et al., 2025) shows promising performance in multi-objective optimization by combining sequential modeling with structured tree-based molecule representation. Reinforcement learning approaches, e.g., MolDQN (Zhou et al., 2018) and GFlowNets (Bengio et al., 2023; Jain et al., 2023), enable property-driven optimization, often with multi-objective extensions for QED, LogP, and SA. However, these single-agent methods struggle to exactly satisfy multiple numeric constraints, reflecting exploration–exploitation trade-offs.
@@ 4.1 Experimental Setup
Table 1: Overall metrics across methods. Lower is better for error metrics; higher is better for diversity and uniqueness. Best per column in bold; second best underlined.
@@
Table 2: Performance comparison across methods on HOMO/LUMO–constrained molecular generation.
@@ 4.2 Results and Analysis
Protocol.
GRPO is ground-truth–free and reward-based, so performance is not tied to a particular training distribution. To test generalization, we uniformly sample 100 target tuples $(\mathrm{QED},\mathrm{LogP},\mathrm{MW})$ and $\mathrm{HOMO},\mathrm{LUMO}$ across admissible ranges. For each tuple and each baseline we run 10 independent trials under the same compute budget and report the best-of-10. Across settings, our normalized total error (NTE) decreases monotonically with hop count shown in both Table 1 and Table 2.
Main Results.
@@ 4.3 Ablation Study
Interpretation.
We ablate three design choices on a held-out set: (i) Stage I without retrieval (baseline), (ii) Stage 1 with retrieval, and (iii) Stage I with retrieval followed by a fragment-level optimizer using 1/2/3 hops. We report per-property errors (QED, logP, MW) and the normalized total error ($e_{\mathrm{norm}}=\lvert\Delta\mathrm{QED}\rvert+\lvert\Delta\log P\rvert/10+\lvert\Delta\mathrm{MW}\rvert/700$) in Table 3. For visualization, we plot the drop percentage relative to the no-retrieval baseline,
@@
Effect of Retrieval.
Adding retrieval already yields consistent gains: the normalized total error drops by 13.7% ($0.307\to 0.265$), driven primarily by improvements in logP ($\mathbf{20.7\%}$ drop) and MW ($7.8\%$ drop). Retrieval also gives the best stand-alone QED error among non-optimized variants ($\mathbf{0.098}$, $\mathbf{11.7\%}$ drop).
@@
Effect of the Fragment-level Optimizer.
Introducing the optimizer produces the largest improvements, especially on MW. Moving from retrieval-only to 1/2/3 hops reduces MW error from $\sim 63$ to $\sim 10$ ($\mathbf{84.9\%}$ drop vs. baseline), and steadily improves logP (drops of $56.4\%$, $65.1\%$, and $\mathbf{70.7\%}$). The overall normalized error decreases monotonically with more hops: $0.187$ (1-hop, $39.1\%$ drop), $0.160$ (2-hop, $47.9\%$ drop), and $\mathbf{0.146}$ (3-hop, $\mathbf{52.4\%}$ drop). QED exhibits a small regression at 1-hop (as expected when trading off multi-objective targets), but recovers by 3-hop to a $7.0\%$ drop versus baseline.
Takeaway.
Retrieval is a strong enabler, and the fragment-level optimizer is essential for precise multi-property alignment, culminating in the best overall performance with the 3-hop setting.